OpenStudy (anonymous):

please how work! Solve using the multiplication principle. -6n < 30

OpenStudy (anonymous):

Just like any other equation. Only if you take the negative sign over you swing around the <

OpenStudy (anonymous):

so its -6n<30=-6n

OpenStudy (anonymous):

-6n<30 -6n=30 -6n/6n= 30/-6 n>-5